OpenStudy (anonymous):

Solve for x: |2x + 12| = 18

OpenStudy (anonymous):

a. x = 3, x = −3 b. x = 3, x = −15 c. x = 15, x = −15 d. x = −3, x = 15

OpenStudy (jadatamiamaraj):

its d

OpenStudy (anonymous):

yep D

OpenStudy (anonymous):

Wait no A

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

| 2x + 12 | = 18 2x + 12 = 18 -(2x + 12) = 18 2x = 18 - 12 -2x - 12 = 18 2x = 6 -2x = 18 + 12 x = 3 -2x = 30 x = -15

OpenStudy (texaschic101):

questions ??

OpenStudy (radar):

Glad someone steered them in correct direction.

OpenStudy (texaschic101):

lol.....I probably shouldn't have given the answer.....just the equations to solve it.

OpenStudy (radar):

That is much better than giving "wrong" answers lol